# Thermoelectric processes of quantum normal-superconductor interfaces

### v1

@@ I Introduction
It is instructive to recall that the existence of thermoelectricity in bulk superconductors was firstly predicted in a series of seminal works by Ginzburg [48, 49, 50] which have motivated experimental research [51, 52]. Thermoelectric [53] and thermal [54, 55, 56] transport in Josephson junctions induced by quasiparticle-Cooper pair coupling have been measured [57] and exploited in a number of coherent caloritronic devices [58, 59]. Recently, strong thermoelectrical effects have been reported in superconducting tunnel junctions done with ferromagnetic superconductors [28, 60, 61, 62], in SIS (I being an insulating barrier) junctions with asymmetric gap configurations [63, 64], magnetic molecules [65], or even Abrikosov vortices [66]. Intriguingly, phase-dependent thermoelectric effects have also been discussed in Andreev interferometers, where good agreement between semiclassical theory in diffusive regimes  [67, 68, 69, 70] or scattering theory for mesoscopic diffusive [29] or chaotic cavities [71, 72] have been contrasted with experiments [73, 74, 75]. In this review, we mainly concentrate on hybrid configurations between normal, topological materials and superconductors, and hence we will not discuss all-superconductor or superconductor-ferromagnetic interfaces, referring the reader to specific reviews [76, 77, 33].
@@ II Physical processes
Thermoelectricity in mesoscopic systems has been reviewed in detail in Ref. 25 and the expected consequences of the different symmetry classes on the linear thermoelectric coefficients have been discussed in Ref. 71. In the presence of one or more superconducting terminals, we refer the reader to the well-known literature [29, 78, 79]. We now briefly review the basic results.
@@ II.1 Charge and heat transport
A general description of transport is given in terms of scattering theory [78, 83]. The full scattering matrix allows us to calculate transport properties. Specifically, it introduces the scattering probabilities
@@
where the sum is expressed over all $m$($m^{\prime}$)-modes coming from the the $l$ ($l^{\prime}$) terminal. For simplicity, we limit our discussion to the current in the normal leads, which is given by the sum of the contributions of all the propagating states in the terminal. [^1] In superconducting terminals the charge current is the sum of the current from the propagating states plus the dissipationless supercurrent from the condensate. In this case, the current carried by $\alpha$-type quasiparticles entering the scatterer from terminal $l$ at a given energy $E$ is
${\cal I}_{l}^{\alpha}(E)=\frac{1}{h}\sum_{l^{\prime},\alpha^{\prime}}\left[{\cal N}_{l}^{\alpha}\delta_{ll^{\prime}}\delta_{\alpha\alpha^{\prime}}-{\cal T}_{ll^{\prime}}^{\alpha\alpha^{\prime}}(E)\right]f_{l^{\prime}}^{\alpha^{\prime}}(E),$ (4)
With the index $l^{\prime}$ running over all terminals and the sum over $\alpha^{\prime}$ including both electron and hole states.  [^2] Note that the sum of $l^{\prime}$ is done on all the terminals which have asymptotically propagating modes at energy $E$ in the terminal. This means that superconducting terminals need to be properly included in the sum only for $E\geq\Delta$. The Fermi distributions are $f_{l}^{\alpha}(E)=\{1+\exp[E-s_{\alpha}\mu_{l})/k_{\text{B}}T_{l}]\}^{-1}$, where $\mu_{l}$ and $T_{l}$ are, respectively, the chemical potential and the temperature of terminal $l$. In this simplified treatment, we assume that all the superconducting terminals are kept at the same chemical potential, which is taken as the reference bias, i.e., $\mu_{S}=0$. Note that with this unified notation ${\cal T}_{ll}^{\alpha\beta}$, with $l=l^{\prime}$ ($l\neq l^{\prime}$), represents Andreev reflection (transmission) probabilities when $\alpha\neq\beta$, and normal reflections (transmissions) for $\alpha=\beta$. Summing over the energies in Eq. (4), one can easily compute the total particle and heat currents [25]
@@ II.2 Linear response
In order to discuss the charge, heat and thermoelectrical transport it is frequently convenient to compute the linear response coefficients. In the linear regime of the normal terminal biases $\delta V_{l}=(\mu_{l}-\mu_{S})/e$ and small temperature gradients $\delta T_{ll^{\prime}}\ll T$, with $T$ the operating temperature, we can write the Onsager equations [89, 25, 90]
@@
where the Onsager coefficients $\mathbf{L}_{ll^{\prime}}$ are expressed in terms of the multiterminal affinities $\delta\mu_{l}/T$ and $\delta T_{ll^{\prime}}/T^{2}$. The linear coefficients reduce to
$\mathbf{L}_{ll^{\prime}}=T\int\frac{\mathrm{d}E}{h}\left[-\frac{\partial f(E)}{\partial E}\right]\begin{pmatrix}e^{2}\ell^{+}_{ll^{\prime}}(E)&eE\ell^{+}_{ll^{\prime}}(E)\\ eE\ell^{-}_{ll^{\prime}}(E)&E^{2}\ell^{-}_{ll^{\prime}}(E)\\ \end{pmatrix},$ (8)
@@
The sign change on ${\cal T}^{he}_{ll^{\prime}}(E)$ is due to the fact that holes contribute with an opposite sign to the charge but have the same contribution as electrons to the heat or energy flow. The Onsager coefficients can be directly connected to the transport coefficients and the most common thermoelectrical figure of merit as reported in Ref. 25. For example, for energy independent coefficients, i.e., $\ell^{\pm}_{ll^{\prime}}(E)\equiv\ell^{\pm}_{ll^{\prime}}$, the multiterminal charge conductance becomes $L^{IV}_{ll^{\prime}}=G_{0}T\ell_{ll^{\prime}}^{+}$ where $G_{0}=e^{2}/h$ is the conductance quantum. Analogously, the thermal term becomes $L^{JT}_{ll^{\prime}}=(\pi^{2}/3)(k_{B}^{2}/e^{2})G_{0}T\ell^{-}_{ll^{\prime}}$. In the absence of superconducting leads and using the Sommerfield expansion in the low temperature limit, the multiterminal Wiedemann-Franz law $L^{JT}_{ll^{\prime}}/L^{IV}_{ll^{\prime}}=(\pi^{2}/3)(k_{B}^{2}/e^{2})T$ is recovered. However, this law is violated when superconductors are present. In fact, when Andreev contributions are present, one generally finds that $\ell^{-}_{ik}/\ell^{+}_{ik}\neq 1$.
@@
the linear thermoelectrical coefficients $L^{IT}$ and $L^{JV}$ are necessarily zero. Equation (10) represents a symmetry between positive and negative energy states. It is, however, different from the particle-hole symmetry ${\cal T}^{\alpha\beta}_{ll^{\prime}}(E)={\cal T}^{-\alpha-\beta}_{ll^{\prime}}(-E)$ that stems from the Nambu representation of the BdG Hamiltonian, Eq. (2). While particle-hole symmetry is always true for any transmission probabilities, the symmetry in Eq. (10) can be broken in multiterminal configurations or by additional degrees of freedom (e.g., spin, sublattice), enabling finite thermoelectrical coefficients. Examples of broken EIS are shown below. However, it is important to note that the direct connection between thermoelectric effects and breaking the symmetry in Eq. (10) applies only in the linear regime. In the nonlinear regime or in the presence of interactions it is not easy to find general implications unless following a perturbative approach [91].
@@ II.3 Quasiparticle tunneling
The main contribution of the superconducting terminals tunnel coupled with normal metals is encoded in the transmission probabilities ${\cal T}^{ee}(E)$. Those quantities, for a tunneling process with a superconducting interface, would present a gapped nature with a combined peaked densities of states
@@
Historically tunnel junctions were shown to generate small thermoelectric effects in connection to charge imbalance effects [93, 94, 95]. However, the main research was initially focused on the nonlinear cooling mechanisms for NIS [96, 97, 98] and SINIS configurations [99]. At the same time, the unique capability to actively control heat fluxes in superconducting tunnel junctions led to the discovery of novel heat rectification properties in NIS [100, 101] that could even be controlled by phase modulation [102, 103]. The experimental combination of such effects with semiconducting materials further expands the applicative possibilities in thermometry and refrigeration [104, 105]. Those advancements show the possibility to successfully apply Peltier cooling to advanced quantum technologies such as QED resonator and, more generally, quantum computing platforms [106, 107]. In the tunneling regime, with at least one superconducting terminal and one gapped terminal, a bipolar thermoelectrical effect induced by spontaneous particle-hole symmetry breaking has been recently predicted [63, 108, 109] and experimentally demonstrated [64, 110]. Even if this intriguing physics depends on interaction effects [111, 112, 113] it is essentially due to a nonlinear temperature difference [114, 115, 116]. More recently, discussions on the thermoelectric properties for the junctions beyond the tunneling (opaque) barrier limit have also been reported [117, 118].
@@ II.4 Andreev reflection
Unlike conventional even-frequency superconductivity, odd-frequency pairing involves a pair amplitude that is antisymmetric under time (or frequency) exchange, allowing spin-triplet $s$-wave states compatible with Fermi statistics. Breaking spin-rotation symmetry at superconducting junction induces spin-triplet Cooper pairs that enable Andreev processes with a significant thermoelectric contribution [46, 47]. In such scenarios, a finite thermopower arising from Andreev-like processes is only possible in the presence of odd-frequency pairing [46]. This principle holds generally for superconducting systems. Importantly, a finite Andreev thermopower requires coexistence of even- and odd-frequency components, a typical feature of proximity-induced unconventional pairing.
@@ III Quantum dots and single-electron transistors
Figure 2: (a) Scheme of a two-terminal thermoelectric engine with a proximized quantum dot. (b) Superconducting density of states for $\Delta=2k_{\text{B}}T_{S}$, and (c) generated current for different gaps in the quasiparticle regime, for $U=2k_{\text{B}}T_{S}$ and $T_{N}=2T_{S}$. The gray line in (e), corresponding to the all-normal case $\nu_{S}=\nu_{N}$, is shown for reference. The tunneling rate through the normal contact is $\Gamma_{N}=2\pi|\tau_{N}|^{2}\nu_{N}/\hbar$. (d) Nonlocal three-terminal heat engine generating current between terminals $N$ and $S$ from the heat injected from $H$ via the capacitive coupling of two quantum dots. (e) Cooper pair splitter serving as a nonlocal thermoelectric and refrigerator.
@@
with the energy of the level $\varepsilon_{\sigma}$, the intradot Coulomb interaction $U$, and the number operators $\hat{n}_{\sigma}=\hat{d}_{\uparrow}^{\dagger}\hat{d}_{\downarrow}$, we get the Hamiltonian of the proximized quantum dot. This Hamiltonian can be diagonalized, giving the even-parity ABS
|±$\left|\pm\right\rangle=\mathcal{N}_{\pm}^{-1}(A_{\pm}\left|0\right\rangle-\Gamma_{S}\left|2\right\rangle),$ (14)
[…]
@@
where $\tilde{\varepsilon}=(\sum_{\sigma}\varepsilon_{\sigma}+U)/2$. These superpositions of states with 0 and 2 electrons mediate the exchange of Cooper pairs with the superconductor which can be controlled by gate voltages applied to the quantum dot. The odd states $|\sigma\rangle$, with eigenenergies $E_{\sigma}=\varepsilon_{\sigma}$, are not coupled. Note that the property $A_{+}\to-A_{-}$ by reversing $\tilde{\varepsilon}\to-\tilde{\varepsilon}$, which can be tuned by the quantum dot gate voltages, defines the particle-hole symmetry point. The consequence of this symmetry is that the energy spectrum of the ABS is symmetric in the energy. So even if the transport properties are clearly affected by the proximity effect [134, 135] this even symmetry of the spectrum may determine the EIS of Eq. (10) suppressing the linear thermoelectrical effects. However, it is important to note that in the nonlinear regime thermoelectricity can emerge [91].

### v2

@@ I Introduction
It is instructive to recall that the existence of thermoelectricity in bulk superconductors was firstly predicted in a series of seminal works by Ginzburg [48, 49, 50], which have motivated experimental research [51, 52]. Thermoelectric [53] and thermal [54, 55, 56] transport in Josephson junctions induced by quasiparticle-Cooper pair coupling have been measured [57] and exploited in a number of coherent caloritronic devices [58, 59]. Recently, strong thermoelectrical effects have been reported in superconducting tunnel junctions done with ferromagnetic superconductors [28, 60, 61, 62], in SIS (I being an insulating barrier) junctions with asymmetric gap configurations [63, 64], magnetic molecules [65], or even Abrikosov vortices [66]. Intriguingly, phase-dependent thermoelectric effects have also been discussed in Andreev interferometers, where good agreement between semiclassical theory in diffusive regimes  [67, 68, 69, 70] or scattering theory for mesoscopic diffusive [29] or chaotic cavities [71, 72] have been contrasted with experiments [73, 74, 75]. In this review, we mainly concentrate on hybrid configurations between normal, topological materials and superconductors, and hence we will not discuss all-superconductor or superconductor-ferromagnetic interfaces, referring the reader to specific reviews [76, 77, 33].
@@ II Physical processes
Thermoelectricity in mesoscopic systems has been reviewed in detail in Ref. 25 and the expected consequences of the different symmetry classes on the linear thermoelectric coefficients have been discussed in Ref. 78. In the presence of one or more superconducting terminals, we refer the reader to the well-known literature [29, 79, 80]. We now briefly review the basic results.
@@ II.1 Charge and heat transport
A general description of transport explicitly respecting the first two laws of thermodynamics [25] is given in terms of scattering theory [79, 84]. The full scattering matrix allows us to calculate transport properties. Specifically, it introduces the scattering probabilities
@@
where the sum is expressed over all mm($m^{\prime}$)-modes coming from the the $l$ ($l^{\prime}$) terminal. For simplicity, we limit our discussion to the current in the normal leads, which is given by the sum of the contributions of all the propagating states in the terminal. [^1] In superconducting terminals, the charge current is the sum of the current from the propagating states plus the dissipationless supercurrent from the condensate. In this case, the current carried by $\alpha$-type quasiparticles entering the scatterer from terminal $l$ at a given energy $E$ is
${\cal I}_{l}^{\alpha}(E)=\frac{1}{2h}\sum_{l^{\prime},\alpha^{\prime}}\left[{\cal N}_{l}^{\alpha}\delta_{ll^{\prime}}\delta_{\alpha\alpha^{\prime}}-{\cal T}_{ll^{\prime}}^{\alpha\alpha^{\prime}}(E)\right]f_{l^{\prime}}^{\alpha^{\prime}}(E),$ (4)
where ${\cal N}_{l}^{\alpha}$ being the number of channels in lead $l$, with the index $l^{\prime}$ running over all terminals and the sum over $\alpha^{\prime}$ including both electron and hole states.  [^2] Note that the sum of $l^{\prime}$ is done on all the terminals which have asymptotically propagating modes at energy $E$ in the terminal. This means that superconducting terminals need to be properly included in the sum only for $|E|\geq\Delta$. The Fermi distributions are $f_{l}^{\alpha}(E)=\{1+\exp[(E-s_{\alpha}\mu_{l})/k_{\text{B}}T_{l}]\}^{-1}$, where $\mu_{l}$ and $T_{l}$ are, respectively, the chemical potential and the temperature of terminal $l$. In this simplified treatment, we assume that all the superconducting terminals are kept at the same chemical potential, which is taken as the reference bias, i.e., $\mu_{S}=0$. Note that with this unified notation ${\cal T}_{l^{\prime}l}^{\alpha\beta}$, with $l=l^{\prime}$ ($l\neq l^{\prime}$), represents (crossed) Andreev reflection probabilities when $\alpha\neq\beta$, and normal reflections (transmissions) for $\alpha=\beta$ [^3] The factor $1/2$ in the operator of Eq. (4) comes from the adopted convention to consider states in the Nambu states which extend in energy from $E\in(-\infty,+\infty)$. This means that all the energy integrals, if not differently indicated, have to be defined over that domain. This representation is fully equivalent to the convention restricting only to states in the positive energy domain, $E>0$. The bilateral convention we adopted for specific computations, especially when small temperature differences are considered, such as for linear thermoelectric properties, is more convenient to identify cancellations between terms. Summing over the energies in Eq. (4), one can easily compute the total particle and heat currents [25]
@@
The currents can also be calculated using standard many-body techniques such as Schwinger-Keldysh non-equilibrium Green’s functions based on microscopic Hamiltonians for the normal and superconducting leads (${\cal H}_{l}$), the mesoscopic region (${\cal H}_{\rm M}$), and tunnel-contacts between them. The precise mapping between the scattering theory quantities and the Green function formalism has been discussed in the literature [90, 91, 39]. In particular, the transmission functions can be expressed in terms of the mesoscopic region retarded/advanced Green function connected to all the leads,
${\bf G}^{r/a}_{\rm M}(E)=\lim\limits_{\eta\to 0^{+}}[(E\pm i\eta)\mathbb{1}-{\cal H}_{\rm M}-\bm{\Sigma}_{\text{leads}}]^{-1},$ (7)
with $\mathbb{1}$ the identity matrix and where $\bm{\Sigma_{\text{leads}}}$ is the self-energy taking into account the effect of the leads on the mesoscopic region. The self-energy due to normal-state leads is diagonal in Nambu space and we can define the hybridization matrices $\bm{\Gamma}_{l}^{\alpha}(E)=2\mathrm{Im}\{\bm{\Sigma}_{l}^{\alpha}(E)\}=2\pi\bm{\tau}_{{\rm M},l}\bm{\rho}^{\alpha}_{l}(E)\bm{\tau}_{{\rm M},l}^{\dagger}$, with the matrices $\bm{\rho}^{\alpha}_{l}(E)$ and $\bm{\tau}_{{\rm M},l}$ respectively representing the leads electron or hole density of states and tunnel amplitudes. The resulting expressions for the transmission coefficients between normal leads in terms of the Green functions are
${\cal T}_{ll^{\prime}}^{\alpha\beta}(E)={}$ $\text{Tr}\left[\bm{\Gamma}_{l}^{\alpha}(E){\bf G}^{r,\alpha\beta}_{\rm M}(E)\bm{\Gamma}_{l^{\prime}}^{\beta}(E){\bf G}^{a,\beta\alpha}_{\rm M}(E)\right],\,\l \neq l^{\prime},$
${\cal T}_{ll}^{\alpha\overline{\alpha}}(E)={}$ $\text{Tr}\left[\bm{\Gamma}_{l}^{\alpha}(E){\bf G}^{r,\alpha\overline{\alpha}}_{\rm M}(E)\bm{\Gamma}_{l}^{\overline{\alpha}}(E){\bf G}^{a,\overline{\alpha}\alpha}_{\rm M}(E)\right],$ (8)
where $\alpha,\beta=e,h$ and $\overline{\alpha}=h,e$ when $\alpha=e,h$. The first line represents the transmission probability between different leads, while the second one corresponds to Andreev reflection processes. Equation (4) is recovered by using the property $\sum_{\beta}{\cal T}_{ll}^{\alpha\beta}(E)={\cal N}_{l}^{\alpha}$. It is important to note that Green function approaches could be very convenient to investigate interaction effects. Typically, this is done following diagrammatic methods [92]. However, discussions using that formalism are typically less transparent than the scattering approach, which instead allows us to visualize the physical processes more clearly. Finally, for linear properties, which we will mainly consider in this review, the two approaches are fully equivalent.
@@ II.2 Linear response
To discuss the charge, heat, and thermoelectrical transport, it is frequently convenient to compute the linear response coefficients. In the linear regime of the normal terminal biases $\delta V_{l}=(\mu_{l}-\mu_{S})/e$ and small temperature gradients $\delta T_{ll^{\prime}}\ll T$, with $T$ the operating temperature, we can write the Onsager equations [78, 25, 93]
@@
where the Onsager coefficients $\mathbf{L}_{ll^{\prime}}$ are expressed in terms of the multiterminal affinities $\delta V_{l}/T$ and $\delta T_{ll^{\prime}}/T^{2}$. The linear coefficients reduce to
$\mathbf{L}_{ll^{\prime}}=T\int\frac{\mathrm{d}E}{2h}\left[-\frac{\partial f(E)}{\partial E}\right]\begin{pmatrix}e^{2}\ell^{+}_{ll^{\prime}}(E)&eE\ell^{+}_{ll^{\prime}}(E)\\ eE\ell^{-}_{ll^{\prime}}(E)&E^{2}\ell^{-}_{ll^{\prime}}(E)\\ \end{pmatrix}$, (10)
@@
The sign change on ${\cal T}^{he}_{ll^{\prime}}(E)$ is because holes contribute with an opposite sign to the charge but have the same contribution as electrons to the heat or energy flow. The Onsager coefficients can be directly connected to the transport coefficients and the most common thermoelectrical figure of merit as reported in Ref. 25. For example, for energy independent coefficients, i.e., $\ell^{\pm}_{ll^{\prime}}(E)\equiv\ell^{\pm}_{ll^{\prime}}$, the multiterminal charge conductance becomes $L^{IV}_{ll^{\prime}}=G_{0}T\ell_{ll^{\prime}}^{+}$ where $G_{0}=e^{2}/h$ is the conductance quantum. Analogously, the thermal term becomes $L^{JT}_{ll^{\prime}}=(\pi^{2}/3)(k_{B}^{2}/e^{2})G_{0}T\ell^{-}_{ll^{\prime}}$. In the absence of superconducting leads and using the Sommerfield expansion in the low temperature limit, the multiterminal Wiedemann-Franz law $L^{JT}_{ll^{\prime}}/L^{IV}_{ll^{\prime}}=(\pi^{2}/3)(k_{B}^{2}/e^{2})T$ is recovered. However, this law is violated when superconductors are present. In fact, when Andreev contributions are present, one generally finds that $\ell^{-}_{ik}/\ell^{+}_{ik}\neq 1$.
@@
the linear thermoelectrical coefficients $L^{IT}$ and $L^{JV}$ are necessarily zero. Equation (12) represents a symmetry between positive and negative energy states. It is, however, different from the particle-hole symmetry ${\cal T}^{\alpha\beta}_{ll^{\prime}}(E)={\cal T}^{-\alpha-\beta}_{ll^{\prime}}(-E)$ that stems from the Nambu representation of the BdG Hamiltonian, cf. Eq. (2). While particle-hole symmetry is always true for any transmission probabilities, the symmetry in Eq. (12) is not generic and can be broken, e.g., in multiterminal configurations or by additional degrees of freedom (e.g., spin, sublattice), enabling finite thermoelectrical coefficients. Examples of broken EIS are shown below. However, it is important to note that the direct connection between thermoelectric effects and breaking the symmetry in Eq. (12) applies only in the linear regime. In the nonlinear regime or in the presence of interactions it is not easy to find general implications unless following a perturbative approach [94].
@@ II.3 Quasiparticle tunneling
The main contribution of the superconducting terminals tunnel coupled with normal metals is encoded in the transmission probabilities ${\cal T}^{ee}(E)$. Those quantities, for a tunneling process with a superconducting interface, would present a gapped nature with a combined peaked density of states
@@
Historically, tunnel junctions were shown to generate small thermoelectric effects in connection to charge imbalance effects [96, 97, 98]. However, the main research was initially focused on the nonlinear cooling mechanisms for NIS [99, 100, 101] and SINIS configurations [102]. At the same time, the unique capability to actively control heat fluxes in superconducting tunnel junctions led to the discovery of novel heat rectification properties in NIS [103, 104] that could even be controlled by phase modulation [105, 106]. The experimental combination of such effects with semiconducting materials further expands the applicative possibilities in thermometry and refrigeration [107, 108]. Those advancements show the possibility to successfully apply Peltier cooling to advanced quantum technologies such as QED resonators and, more generally, quantum computing platforms [109, 110]. In the tunneling regime, with at least one superconducting terminal and one gapped terminal, a bipolar thermoelectrical effect induced by spontaneous particle-hole symmetry breaking has been recently predicted [63, 111, 112] and experimentally demonstrated [64, 113]. Even if this intriguing physics depends on interaction effects [114, 115, 116] it is essentially due to a nonlinear temperature difference [117, 118, 119]. More recently, discussions on the thermoelectric properties for the junctions beyond the tunneling (opaque) barrier limit [120, 121] and including thermal models accounting for nonlinearities [122] have also been reported.
@@ II.4 Andreev reflection
Unlike conventional even-frequency superconductivity, odd-frequency pairing involves a pair amplitude that is antisymmetric under time (or frequency) exchange, allowing spin-triplet $s$-wave states compatible with Fermi statistics. Breaking spin-rotation symmetry at a superconducting junction induces spin-triplet Cooper pairs that enable Andreev processes with a significant thermoelectric contribution [46, 47]. In such scenarios, a finite thermopower arising from Andreev-like processes is only possible in the presence of odd-frequency pairing [46]. This principle holds generally for superconducting systems. Importantly, a finite Andreev thermopower requires coexistence of even- and odd-frequency components, a typical feature of proximity-induced unconventional pairing.
Finally, in superconducting heterostructures the Andreev approximation assumes that the chemical potential is much larger than the energies involved ($\mu\gg E,\Delta$). This means that electrons and holes in Andreev reflections have nearly the same momentum and follow the same path—called retro-reflection. In Sec. V, we examine what happens in Dirac materials, where this approximation breaks down because of their linear energy-momentum relation.
@@ III Quantum dots and single-electron transistors
Figure 2: (a) Scheme of a two-terminal thermoelectric engine with a proximized quantum dot. (b) Superconducting density of states for $\Delta=2k_{\text{B}}T_{S}$, and (c) generated current for different gaps in the quasiparticle regime, for $U=2k_{\text{B}}T_{S}$ and $T_{N}=2T_{S}$. The gray line in (c), corresponding to the all-normal case $\nu_{S}=\nu_{N}$, is shown for reference. The tunneling rate through the normal contact is $\Gamma_{N}=2\pi|\tau_{N}|^{2}\nu_{N}/\hbar$. (d) Nonlocal three-terminal heat engine generating current between terminals $N$ and $S$ from the heat injected from $H$ via the capacitive coupling of two quantum dots. (e) Cooper pair splitter serving as a nonlocal thermoelectric and refrigerator.
@@
with the energy of the level $\varepsilon_{\sigma}$, the intradot Coulomb interaction $U$, and the number operators $\hat{n}_{\sigma}=\hat{d}_{\sigma}^{\dagger}\hat{d}_{\sigma}$, we get the Hamiltonian of the proximized quantum dot. The interaction term lifts the single particle description of the scattering theory, so many body descriptions are required e.g., Keldysh nonequilibrium Green functions [92] or master equations [138]. This Hamiltonian can be diagonalized, giving the even-parity ABS
$\left|\pm\right\rangle=\mathcal{N}_{\pm}^{-1}(A_{\pm}\left|0\right\rangle-\Gamma_{S}\left|2\right\rangle),$ (16)
where $\mathcal{N}_{\pm}$ is imposed by the normalization of the $\left|\pm\right\rangle$ states, with energies
@@
where $\tilde{\varepsilon}=(\sum_{\sigma}\varepsilon_{\sigma}+U)/2$. These superpositions of states with 0 and 2 electrons mediate the exchange of Cooper pairs with the superconductor which can be controlled by gate voltages applied to the quantum dot. The odd states $|\sigma\rangle$, with eigenenergies $E_{\sigma}=\varepsilon_{\sigma}$, are not coupled. Note that the property $A_{+}\to-A_{-}$ by reversing $\tilde{\varepsilon}\to-\tilde{\varepsilon}$, which can be tuned by the quantum dot gate voltages, defines the particle-hole symmetry point. The consequence of this symmetry is that the energy spectrum of the ABS is symmetric in energy. So even if the transport properties are clearly affected by the proximity effect [139, 135] this even symmetry of the spectrum may determine the EIS of Eq. (12), suppressing the linear thermoelectrical effects. However, it is important to note that in the nonlinear regime thermoelectricity can emerge [94, 140].